​(Inflation and project cash​ flows)  If the price of a gallon of regular gasoline is $2.44 and the anticipated rate of inflation in energy prices is such that the cost of gasoline is expected to rise by 8 percent per​ year, what is the expected price per gallon in 9 ​years? The expected price per gallon of gas in 9 years is ​$nothing. ​(Round to two decimal​ places.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=2.44*(1.08)^9

=2.44*1.999004627

=$4.88(Approx).